Suppose the mean number of days to germination of a variety of seeds is 34, with a standard deviation of 4.3 days. What is the probability that the mean germination time of a sample of 250 seeds will be within 0.5 days of the population mean?



1
Expert's answer
2022-03-18T15:20:50-0400

We have a normal distribution, "\mu=34, \sigma=4.3, n=250."

Let's convert it to the standard normal distribution, "z=\cfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}" ;

"z_1=\cfrac{33.5-34}{4.3/\sqrt{250}}=-1.84, \\ z_2=\cfrac{34.5-34}{4.3/\sqrt{250}}=1.84,"

"P(33.5<\bar{X}<34.5)=\\
=P(-1.84<Z<1.84)=\\
=P(Z<1.84)-P(Z<-1.84)="

"=0.9671-0.0329=0.9342" (from z-table).